### v1

Abstract
We investigate the energy efficiency of a library designed for parallel computations with sparse matrices. The library leverages high-performance, energy-efficient Graphics Processing Unit (GPU) accelerators to enable large-scale scientific applications. Our primary development objective was to maximize parallel performance and scalability in solving sparse linear systems whose dimensions far exceed the memory capacity of a single node.
To this end, we devised methods that expose a high degree of parallelism while optimizing algorithmic implementations for efficient multi-GPU usage. Previous work has already demonstrated the library’s performance efficiency on large-scale systems comprising thousands of NVIDIA GPUs, achieving improvements over state-of-the-art solutions.
In this paper, we extend those results by providing energy profiles that address the growing sustainability requirements of modern HPC platforms. We present our methodology and tools for accurate runtime energy measurements of the library’s core components and discuss the findings. Our results confirm that optimizing GPU computations and minimizing data movement across memory and computing nodes reduces both time-to-solution and energy consumption. Moreover, we show that the library delivers substantial advantages over comparable software frameworks on standard benchmarks.
1 Introduction
Sustainability is becoming a major concern in large-scale scientific computing. While high-performance computing platforms advance in computational and communication capabilities, power constraints pose significant challenges, impacting operational costs and system reliability up to the point of being a pivotal issue for organizations and nations alike. The energy consumption of data centers received considerable attention in the International Energy Agency’s 2025 report. In particular, the rapid expansion of data centers in countries such as China and the United States [^1] In the United States, data centers are expected to account for approximately $11.7\%$ of the country’s total electricity demand by 2030. makes the sector one of the major drivers of electricity demand growth, with substantial implications for national energy landscapes [1]. Achieving optimal efficiency in power usage is essential for enabling sustainable HPC infrastructures by minimizing operational costs [2]. The rise of Green IT, which advocates for environmentally sustainable computing, has intensified efforts to improve the energy efficiency of HPC systems. The Green500 list [3], which ranks the world’s most energy-efficient supercomputers based on their performance in FLOPS per Watt, underscores this growing emphasis on energy-efficient supercomputing, highlighting systems with superior performance-to-power ratios.
Notably, the adoption of Graphics Processing Units (GPUs) boosted energy-efficient computing, particularly in applications that benefit from parallel processing [4, 5, 2]. Compared to Central Processing Units (CPUs), GPUs offer superior computational and energy efficiency, especially for high-throughput, high-latency workloads such as scientific machine learning and image processing. As a result, most modern HPC platforms are hybrid systems made of CPUs and GPUs tightly coupled to each other. The widespread adoption of GPU-accelerated HPC systems is well reflected in the Top500 [6] and Green500 rankings of the world’s fastest and most energy-efficient supercomputers. Specifically, in the most recent Green500 list, all of the top ten supercomputers are equipped with GPUs, eight of which feature NVIDIA chips. So, it is essential to assess the energy consumption of GPU sub-systems in today’s power-constrained computing landscape. To address this challenge, hardware and software strategies have been developed to optimize power usage in supercomputing applications [7, 8].
The primary goal of power management in high-performance computing is to minimize energy consumption while maintaining computational performance within predefined limits. Higher energy efficiency can be achieved by reducing either average power consumption or execution times. Research indicates that source-code transformations and application-specific optimizations can significantly enhance GPU resource utilization, performance, and energy efficiency [5]. Substantial energy savings can be achieved by refining GPU implementations and addressing performance bottlenecks. The authors of [2] highlight that even greater efficiency gains are possible by utilizing optimized numerical libraries developed by experts to leverage hardware features and reduce runtimes.
In this paper, we address the critical issue of energy consumption in the core functionalities of BootCMatchGX, a parallel library for sparse matrix computations. The library is the outcome of a numerical software development project focused on designing and implementing scalable sparse linear solvers and preconditioners for NVIDIA GPU-accelerated supercomputers. It includes comprehensive support for Krylov subspace methods, incorporating Sparse Basic Linear Algebra operations—such as the sparse matrix-vector product (SpMV)—that are specifically optimized for efficiency on heterogeneous clusters. Additionally, it provides essential operations for Algebraic MultiGrid (AMG) preconditioners. We present a methodology for analyzing the energy consumption profiles of both the fundamental operations and the overall linear solver, utilizing software tools that enable access to internal hardware sensors [4]. While previous studies have mainly addressed its performance and scalability, here we complement those results with a detailed energy analysis. Our aim is to provide a comprehensive assessment of both runtime efficiency and energy consumption, which is crucial for sustainable high-performance computing. The main contributions of this paper can be summarized as follows:
  • We present a methodology for fine-grained power measurement based on internal CPU and GPU sensors, integrating the LIKWID toolset with our powerMonitor utility.
  • We provide detailed energy profiles of the library’s fundamental building blocks, including sparse matrix–vector multiplication (SpMV), Conjugate Gradient (CG), and Preconditioned Conjugate Gradient (PCG) solvers, offering a holistic view of performance-to-energy trade-offs.
  • We carry out extensive strong and weak scalability experiments using up to 64 NVIDIA GPUs, and compare BootCMatchGX against state-of-the-art frameworks such as Ginkgo and NVIDIA AmgX.
Our results show that BootCMatchGX consistently achieves lower execution times and reduced dynamic energy consumption compared to Ginkgo. In addition, in the PCG case, BootCMatchGX outperforms NVIDIA AmgX due to the improved convergence properties of its preconditioner and fine-tuning GPU implementation of basic operations. These findings confirm the effectiveness of algorithmic optimizations and communication-reduction strategies for both performance and energy sustainability.
The remainder of this paper is organized as follows. Section 2 collocates this work in the current literature on the energy efficiency aspects of sparse matrix computations on GPU-accelerated systems. Section 3 presents the design and main features of the BootCMatchGX library. Section 4 describes the methodology and tools employed for power and energy measurements. Section 5 reports and compares the experimental results for SpMV, Conjugate Gradient, and Preconditioned Conjugate Gradient computations on multi-GPU clusters. Finally, Section 6 summarizes the main findings and outlines directions for future work.
2 Energy efficiency in sparse matrix computations for GPU-accelerated systems
Sparse matrix computations are the core of many scientific applications, ranging from traditional simulation models based on Partial Differential Equations (PDEs) to more recent scientific machine learning approaches [9, 10]. In particular, iterative Krylov solvers are the methods of choice for solving large, sparse linear systems involving hundreds of billions of equations, problems that can be efficiently handled using current petascale and pre-exascale high-performance computing (HPC) systems. Profiling the energy consumption of such computations, alongside traditional performance metrics, provides a more comprehensive understanding of the overall efficiency of existing software frameworks.
In general, sparse matrix computations exhibit a memory-bound behavior due to their low operational intensity. Consequently, they are more sensitive to memory and network bandwidth limitations than to the floating-point throughput of the underlying architecture. This characteristic presents a significant challenge on heterogeneous systems with deep memory hierarchies, where the energy cost of data movement is often orders of magnitude higher than that of performing a double-precision floating-point operation [11, 12, 13]. This has led to growing interest in analyzing the energy efficiency of sparse matrix computations on GPU-accelerated systems.
Several studies have conducted experiments and comparisons across different implementations of key computational kernels—such as sparse matrix-vector multiplication (SpMV)—on a range of test cases. In [5], the authors compare the performance and energy efficiency of SpMV implementations from cuSPARSE and MAGMA for GPUs, as well as Intel’s MKL for multicore CPUs, on the Swiss supercomputer Piz Daint. Their findings show that optimized GPU code can significantly enhance both the computational and energy efficiency of scientific applications. A comprehensive review of state-of-the-art SpMV implementations on GPUs, including the use of machine learning techniques to select the most efficient method in terms of runtime and energy consumption, is provided in [14].
In [15] authors analyzed the power consumption of GPU-accelerated Generalized Minimal Residual (GMRES) solvers enhanced with preconditioning, mixed-precision iterative refinement, and CPU-focused power-saving techniques such as idle-wait periods and Dynamic Voltage and Frequency Scaling (DVFS). While these methods can reduce energy consumption by $6-10\%$, they are less effective for solvers like CG, where the SpMV dominates and leaves little idle time for the CPU. A more recent study [16] introduced batched sparse and mixed-precision linear algebra interfaces tailored for applications involving many small-scale systems. By grouping operations into batches and exploiting mixed-precision arithmetic, the approach improves GPU utilization and achieves notable reductions in energy-to-solution. Furthermore, [17] presents a detailed analysis of the performance and energy footprint of the CG method combined with Gauss-Seidel-based preconditioners, specifically designed for GPUs, across various GPU architectures.
3 BootCMatchGX library
BootCMatchGX is the latest development in a mathematical software project aimed at designing new methods and efficient implementations of iterative Krylov solvers and algebraic multigrid preconditioners for solving sparse linear systems. It expands and enhances the sequential library BootCMatch [18] and its Nvidia GPU version, BootCMatchG [19, 20]. Its design is driven by the need to scale the library to thousands of GPUs, enabling the solution of systems with many billions of degrees of freedom (DOFs). Scalability and high performance efficiency, specifically the efficient use of GPUs, have been the main guidelines in all the phases of the software development, from design of parallel algorithms to their implementation, as described in [21, 22].
Sparse matrix computations represent a significant computational bottleneck in many scientific applications across various domains. They are ubiquitous in both traditional physics-based modeling and simulation, as well as in data-driven approaches such as variational data assimilation and machine learning. Specifically, preconditioned Krylov methods are the preferred approach for iteratively solving sparse linear systems in very large dimensions, as they preserve the system matrix and are therefore compatible with compressed storage schemes, thus avoiding the complications associated with the typical fill-in phenomenon encountered in direct methods. On the other hand, the low operation intensity, i.e., the low FLOP-to-byte ratio of SpMV, which is the core operation in Krylov methods, makes these methods sub optimal for current supercomputers, both in terms of performance and energy efficiency.
BootCMatchGX is an extensible software library available in source form [^2] BootCMatchGX is available at https://github.com/bootcmatch/BootCMatchGX.. Written in the C programming language, it employs MPI for data communication among parallel tasks and leverages the NVIDIA CUDA framework to exploit the computational power of NVIDIA GPUs. The library provides all essential functionalities for implementing several variants of the well-known Conjugate Gradient (CG) method. These include scalar products of dense vectors (dot), dense vector updates (axpy), norm computations, and SpMV, all in a distributed-memory parallel setting. At the task level, all computations are optimized for efficient execution on NVIDIA GPUs.
Sparse matrices are stored using the Compressed Sparse Row (CSR) format and are distributed across parallel tasks in blocks of contiguous rows. Special care has been taken to handle the challenges of solving systems with more than $4\times 10^{9}$ degrees of freedom (DOFs) on thousands of GPUs. Using 8-byte integers for row and column indices would significantly increase memory requirements. Additionally, on GPUs, 8-byte integers introduce considerable overhead. To avoid this, the library maps global-to-local column indices using a shift mechanism. Specifically, on each GPU, the local column index is computed as the global column index (which exceeds $2^{32}-1$) minus the global index of the first row handled by that GPU. According to this convention, some column indices may temporarily become negative. However, this is only an intermediate state. Before the matrix is used, column indices are compacted and re-numbered so that all operations involve indices starting from zero. The only constraint is that the number of distinct column indices on each GPU must not exceed $2^{32}-1$, as local indices are stored in 4-byte integers. This is generally not a significant limitation, especially for sparse problems arising from PDE discretization. Importantly, there is no restriction on the number of distinct column indices in the global matrix, provided enough GPUs are available to distribute the matrix.
Communication-reduction strategies have been employed throughout the library, from the design of numerical algorithms to their implementation. These include maximizing data reuse at near-thread memory levels, minimizing host-to-GPU (and GPU-to-host) memory transfers, and overlapping GPU-level computation with inter-node communication wherever possible.
At the solver level, BootCMatchGX includes three distinct variants of the PCG method for solving sparse linear systems with symmetric positive-definite (SPD) coefficient matrices. These variants are: the classical algorithm originally introduced by Hestenes and Stiefel [23]; a communication-reduced variant of the flexible Conjugate Gradient method, as proposed in [24]; and the s-step Conjugate Gradient method developed by Chronopoulos and Gear [25]. Further details on these algorithms and the implementation design patterns tailored for multi-GPU clusters can be found in [21, 22].
The development of BootCMatchGX was initially motivated by the goal of providing a robust and scalable Algebraic MultiGrid (AMG) preconditioner—originally proposed in [18]—that offered improved convergence properties and performance efficiency compared to similar methods available in existing libraries, such as NVIDIA AmgX [26]. This AMG preconditioner is based on a coarsening strategy that aggregates degrees of freedom (DOFs) using a maximum-weight matching on a weighted graph derived from the adjacency graph of the system matrix. The coarsening procedure, known as Compatible weighted Matching, has been specifically redesigned to maximize parallelism while preserving, as much as possible, the convergence properties of the preconditioner, as detailed in [21].
In this work, we aim to analyze the energy efficiency profiles of the basic SpMV functionality, as well as of the main solver and preconditioner provided by BootCMatchGX, and compare them with similar functionalities available in state-of-the-art libraries.
4 Energy Consumption Measurements
Measuring power consumption directly using internal or external hardware sensors is widely regarded as the most accurate method for energy assessment [4]. Energy usage can be estimated by periodically sampling sensor readings during an application’s execution, and the total energy consumed is then computed by integrating the power-time curve over the execution interval. Many hardware components include built-in sensors that expose power management interfaces, allowing users to monitor power usage in real time. These internal sensors are convenient, require no additional cost or hardware setup, and support fine-grained, component-level profiling. In contrast, external power measurement devices—while potentially more flexible and accurate—can be costly and impractical for large-scale or distributed systems.
In this study, we rely on internal hardware sensors to monitor CPU and GPU power consumption during execution. Our primary objective is to characterize the energy footprint of BootCMatchGX and other comparable state-of-the-art libraries. Rather than investigating power-saving techniques, such as DVFS, for improving GPU energy efficiency, our focus is on providing a detailed and objective assessment of the energy behavior exhibited by our application software, also in comparison with other solutions.
4.1 CPU and GPU Power
On-chip power sensors integrated into modern CPU and GPU platforms provide high-frequency power measurements that are accessible through a specialized API. A prominent example is Intel’s Running Average Power Limit (RAPL) interface [27]. RAPL is available on Intel multicore CPUs and enables accurate monitoring of energy consumption across multiple components, including CPU cores, DRAM, and integrated GPUs. Energy usage is tracked via 32-bit Model Specific Registers (MSRs), which store cumulative energy readings since the processor was powered on. These counters are typically updated every millisecond. RAPL data can be accessed through a variety of programmatic and command-line tools, such as the Linux sysfs interface, performance monitoring events (perf), or the LIKWID tool suite [28, 29]. LIKWID is a lightweight and user-friendly set of command-line utilities and libraries designed for performance-oriented developers. It supports a range of architectures, including Intel, AMD, ARMv8, and POWER9 processors running Linux.
On the GPU side, NVIDIA GPUs are equipped with on-chip power sensors that expose power measurements via the NVIDIA Management Library (NVML) interface [30]. This API reports the power consumption of the GPU and its associated circuitry in milliwatts. According to the NVML documentation, for Ampere GPUs (excluding GA100) and newer architectures, the reported values represent power averaged over a one-second interval. In contrast, for GA100 and older architectures, the API returns instantaneous power readings.
In this work, we monitor the energy consumption of CPUs and GPUs using LIKWID and NVML, respectively. Specifically, we employ the likwid-perfctr tool from the LIKWID suite, which supports various operational modes. We use it in combination with the LIKWID MarkerAPI, a collection of functions and macros that facilitate the measurement of specific code regions. This approach allows for measuring the energy consumption of each kernel while excluding the power required for input data generation. likwid-perfctr monitors the power consumption of the application from the start to the end of the execution. It generates an output file for each core used by the application, containing information on execution time, energy, and power consumption. We do not use LIKWID for GPU monitoring; instead, we access the device directly through NVML. This approach enables us to reconstruct the GPUs’ power–time curve and more accurately estimate their static power.
For GPU power monitoring, we developed a lightweight tool called powerMonitor [31]. The tool is built on top of GPowerU [32], with both frameworks relying on the NVML interface. GPowerU has been developed within the TEXTAROSSA [7, 8] project [^3] https://www.textarossa.eu, a three-year project co-funded by the European High Performance Computing (EuroHPC) JU [^4] https://eurohpc-ju.europa.eu/. It is a simple tool that measures the power consumption of a CUDA kernel at specific points in the device code and generates a complete power profile.
Figure 1: Execution workflow for CPU and GPU power monitoring.
powerMonitor has been adapted from GPowerU to better meet our specific requirements, providing a simple, yet effective, solution for monitoring the power consumption of all GPUs on a compute node. Indeed, it can be used either as an external tool or integrated directly into the application code. In our experiments, we used it as an external tool. In this case, it must be launched before the application starts and stopped once execution is complete. The complete workflow for CPU and GPU power monitoring is shown in Figure 1. powerMonitor creates an output file for each device found on the node, containing a sequence of power samplings along with the corresponding timestamps. These samples can be used to reconstruct the power timeline of each device during the application’s execution. Figure 2 shows an example of a power–time curve.
4.2 Static and Dynamic Energy
The power consumption of a device can be broadly classified into two categories: static power and dynamic power. Static power refers to the energy consumed by the device simply by being powered on, regardless of whether it is performing any computing intensive operations. Dynamic power, on the other hand, is the additional energy consumed when the device executes an application.
Figure 2: Power–time profile of the SpMV kernel measured within the BootCMatchGX library on a single node equipped with four GPUs. The green and purple markers denote the points at which the GPUs leave and return to the idle state, respectively. These reference points are used to estimate the static power consumption of the GPUs.
Accordingly, the static energy consumed during the execution of an application is given by the product of the device’s static power and the application’s execution time. The dynamic energy is then calculated as the difference between the total energy consumed during execution and the static energy [33, 34]. The total energy consumption ($TE$) of an application can be expressed as:
$TE=TP_{GPU}\times T+TP_{CPU}\times T,$
where $T$ is the execution time, and $TP_{GPU}$ and $TP_{CPU}$ denote the average total power for the GPU and CPU, respectively. Similarly, the static energy ($SE$) and dynamic energy ($DE$) are defined as:
$SE=SP_{GPU}\times T+SP_{CPU}\times T,$
DE=TESE,DE=TE-SE,
where $SP_{GPU}$ and $SP_{CPU}$ represent the static power of the GPU and CPU, respectively. In our analysis, we focus on the dynamic energy consumption, which corresponds to the additional energy required to execute the application. To estimate $DE$, we proceed as follows:
  • For the CPU, the total energy consumption $TE_{CPU}$ and the runtime $T$ are obtained using likwid-perfctr. The static power $SP_{CPU}$ is estimated using likwid-powermeter in stethoscope mode, which allows for measurements of the CPU’s power consumption over a specified time interval. We set this interval to 1 second to sample the CPU’s power usage. Before running the experiments, likwid-powermeter is executed on idle computational nodes to measure $SP_{CPU}$, which is then multiplied by the execution time $T$ to calculate the static energy $SE_{CPU}$.
  • For the GPU, the total energy consumption $TE_{GPU}$ is computed by integrating the power timeline obtained from powerMonitor, which samples power approximately $20$ times per millisecond. Figure 2 illustrates an example of the power timeline recorded during the execution of a target application on a single node with $4$ GPUs. To calculate $SE_{GPU}$, we estimate the static power $SP_{GPU}$ from the data and integrate it over the running time.The green and purple markers in Figure 2 denote the transitions of the GPUs between idle and active states.
Finally, we compute the dynamic energies $DE_{GPU}$ and $DE_{CPU}$, and sum them to obtain the total dynamic energy consumption:
$DE=DE_{GPU}+DE_{CPU}.$
5 Results and Comparisons
In this section, we analyze the performance and energy consumption of BootCMatchGX, comparing it with state-of-the-art libraries. Tests were conducted on a cluster with dual-socket Intel Xeon Gold CPUs (32 cores each) and four NVIDIA A30 GPUs per node, interconnected via HDR InfiniBand; up to 16 nodes (64 GPUs) were used. The software stack included CUDA 12.3, Open MPI 4.1.6, and GCC 12.2.1.
We evaluate the SpMV operation, as a fundamental sparse matrix operation, and of the CG solver. BootCMatchGX (v1.1.0) is compared with Ginkgo (v1.9.0) [35, 36] for SpMV and un-preconditioned CG, while preconditioned CG results are also compared with NVIDIA AmgX (v2.4.0) [37, 38]. The AmgX SpMV implementation is not considered, as no official driver is available to test it separately.
Benchmark problems are derived from the 3D Poisson equation with homogeneous Dirichlet boundary conditions, discretized on uniform meshes with 7- and 27-point stencils (the latter as in HPCG [39]). We analyze performance and energy efficiency under both strong and weak scalability conditions. In both cases, the local problem size is initially chosen as the largest that fits into the memory capacity of a single GPU, ensuring full exploitation of device resources and realistic memory-bound conditions. In strong scaling experiments, this global problem (corresponding to the single-GPU memory-saturating size) is kept fixed and partitioned among GPUs as their number increases, leading to a reduced local workload per GPU. In weak scaling experiments, instead, the global problem size grows linearly with the number of GPUs, while the local workload per GPU remains constant at the memory-saturating size. Matrices are partitioned by rows across GPUs, with the 3D domain mapped to a 3D grid of MPI tasks, reproducing realistic communication/computation patterns typical of large-scale PDE-based applications.
5.1 SpMV Operation
We begin by analyzing the performance behavior of the SpMV operation under both strong and weak scalability scenarios.For the single-GPU case, the problem size is set to $405^{3}$ DOFs for the 7-point stencil and $260^{3}$ DOFs for the 27-point stencil, corresponding to the largest dimensions that fully saturate the available GPU memory. All performance measurements reported in the figures represent averages over five independent runs, with each run consisting of 100 repetitions of the SpMV computation.
(a) 7-points stencil matrix with $405^{3}$ DOFs per GPU under weak scalability.
(b) 27-points stencil matrix with $260^{3}$ DOFs per GPU under weak scalability.
(c) 7-points stencil matrix with a total of $405^{3}$ DOFs under strong scalability.
(d) 27-points stencil matrix with a total of $260^{3}$ DOFs under strong scalability.
Figure 3: SpMV execution times under weak and strong scalability scenarios.
Figure 3 shows the execution times of the SpMV computation, comparing the BootCMatchGX and Ginkgo implementations. Across all tested scenarios, BootCMatchGX consistently outperforms Ginkgo, with notably lower execution times. The performance gap is especially pronounced in the weak scalability cases. These results suggest a potential advantage for BootCMatchGX in terms of both performance and energy efficiency for very large-scale computations.
(a) 7-points stencil matrix with $405^{3}$ DOFs per GPU under weak scalability.
(b) 27-points stencil matrix with $260^{3}$ DOFs per GPU under weak scalability.
(c) 7-points stencil matrix with a total of $405^{3}$ DOFs under strong scalability.
(d) 27-points stencil matrix with a total of $260^{3}$ DOFs under strong scalability.
Figure 4: Dynamic energy consumption breakdown of the SpMV computation on GPU and CPU under weak and strong scalability scenarios.
Figure 4 reports the breakdown of dynamic energy consumption for the SpMV computation, with GPU and CPU contributions represented as distinct colored segments in each bar, as indicated in the legend. The results demonstrate that the BootCMatchGX implementation is consistently more energy-efficient than Ginkgo under both weak and strong scalability conditions. In all cases, its dynamic energy consumption is approximately half that of Ginkgo. The data also reveal that the CPU contribution to the overall energy consumption is negligible compared to that of the GPU. This outcome is expected, as the bulk of the computational workload is offloaded to the GPU, while the CPU is primarily responsible for inter-process communication—a task that requires limited computational effort and therefore incurs only a marginal energy cost.
(a) 7-points stencil matrix with $405^{3}$ DOFs per GPU under weak scalability.
(b) 27-points stencil matrix with $260^{3}$ DOFs per GPU under weak scalability.
(c) 7-points stencil matrix with a total of $405^{3}$ DOFs under strong scalability.
(d) 27-points stencil matrix with a total of $260^{3}$ DOFs under strong scalability.
Figure 5: GPU power peak of the SpMV computation under weak and strong scalability scenarios.
Figure 5 shows the GPU power peaks recorded during the execution of the SpMV computation. As the charts clearly illustrate, BootCMatchGX consistently exhibits lower GPU power peaks than Ginkgo, indicating a more efficient use of GPU resources by maintaining a steady workload and avoiding sudden power spikes. As expected, under strong scalability conditions, the power peaks decrease with an increasing number of GPUs, due to the reduced workload per GPU.
(a) 7-points stencil matrix with $405^{3}$ DOFs under weak scalability.
(b) 27-points stencil matrix with $260^{3}$ DOFs under weak scalability.
(c) 7-points stencil matrix with a total of $405^{3}$ DOFs under strong scalability.
(d) 27-points stencil matrix with a total of $260^{3}$ DOFs under strong scalability.
Figure 6: Dynamic energy consumption per DOF breakdown of the SpMV computation under weak and strong scalability scenarios.
Finally, Figure 6 presents the dynamic energy consumption per DOFs. As anticipated from previous results, BootCMatchGX demonstrates significantly higher energy efficiency than Ginkgo—approximately twice as efficient. In both implementations, energy efficiency remains nearly constant with increasing numbers of GPUs under weak scalability, confirming the good scalability of both solutions. Under strong scalability, as expected, efficiency declines as the number of GPUs increases, due to the reduction of the computational workload per GPU.
Overall, our analysis shows that the BootCMatchGX implementation of SpMV outperforms Ginkgo in both execution time and energy consumption. Its improved performance directly contributes to a lower energy footprint. Moreover, the consistently lower GPU power peaks observed for BootCMatchGX (see Figure 5) suggest a more effective utilization of GPU resources, further reinforcing its energy-efficiency features.
(a) 7-points stencil matrix with $408^{3}$ DOFs per GPU under weak scalability.
(b) 27-points stencil matrix with $265^{3}$ DOFs per GPU under weak scalability.
(c) 7-points stencil matrix with a total of $408^{3}$ DOFs under strong scalability.
(d) 27-points stencil matrix with a total of $265^{3}$ DOFs under strong scalability.
Figure 7: Un-preconditioned CG execution times under weak and strong scalability scenarios.
5.2 Un-preconditioned Conjugate Gradient solver
We evaluate the un-preconditioned CG implementations of BootCMatchGX, Ginkgo, and NVIDIA AmgX under both strong and weak scalability. The analysis of the un-preconditioned solver is of particular interest, as it provides a direct insight on the intrinsic computational costs of a CG iteration without the influence of preconditioning. In this setting, the performance can be directly attributed to the efficiency of the SpMV operation—by far the dominant kernel in terms of execution time—as well as to the relative impact of other fundamental operations such as axpy, dot products, and global reductions. This perspective is essential for isolating the contribution of these core building blocks and for assessing their impact independently of any preconditioning strategy.
For the single-GPU case, the problem size is set to $408^{3}$ DOFs with the 7-point stencil and $265^{3}$ DOFs with the 27-point stencil; AmgX is excluded from the 27-point case, as this benchmark is not supported. Reported results are averages over five runs. For our tests, we set the maximum number of iterations to $100$ and the relative residual tolerance to $10^{-16}$. Since our focus is on the cost per iteration rather than on convergence properties, this setup ensures that each implementation performs exactly 100 iterations in all scenarios.
Figure 7 reports the execution times of the CG solver. Across all scenarios, BootCMatchGX consistently outperforms Ginkgo, achieving substantially lower runtimes, particularly under weak scalability. For the 7-point stencil, NVIDIA AmgX also shows superior performance compared to Ginkgo, while delivering results comparable to those of BootCMatchGX. In strong scalability tests, however, the gap narrows as the number of GPUs increases, possibly due to the growing impact of communication and synchronization overheads. Under these conditions, possible advantage of specialized GPU implementations diminishes, and Ginkgo achieves performance comparable to both BootCMatchGX and AmgX. However, we can conclude that both NVIDIA AmgX and BootCMatchGX show clear advantages over Ginkgo in terms of performance, and potentially energy efficiency, for large-scale computations.
(a) 7-points stencil matrix with $408^{3}$ DOFs per GPU under weak scalability.
(b) 27-points stencil matrix with $265^{3}$ DOFs per GPU under weak scalability.
(c) 7-points stencil matrix with a total of $408^{3}$ DOFs under strong scalability.
(d) 27-points stencil matrix with a total of $265^{3}$ DOFs under strong scalability.
Figure 8: Dynamic energy consumption per iteration breakdown of the un-preconditioned CG computation on GPU and CPU under weak and strong scalability scenarios.
Figure 8 breaks down the dynamic energy consumption per iteration for the CG computation, with GPU and CPU contributions shown as colored segments in each bar (see legend). The results indicate that both NVIDIA AmgX and BootCMatchGX are consistently more energy-efficient than Ginkgo under both weak and strong scalability conditions. Their per-iteration dynamic energy consumption is less than half of Ginkgo, while AmgX and BootCMatchGX exhibit comparable overall energy usage. Notably, BootCMatchGX achieves a slight advantage over AmgX in both weak and strong scalability cases.
(a) 7-points stencil matrix with $408^{3}$ DOFs per GPU under weak scalability.
(b) 27-points stencil matrix with $265^{3}$ DOFs per GPU under weak scalability.
(c) 7-points stencil matrix with a total of $408^{3}$ DOFs under strong scalability.
(d) 27-points stencil matrix with a total of $265^{3}$ DOFs under strong scalability.
Figure 9: Dynamic energy consumption per DOF breakdown of the un-preconditioned CG computation on GPU and CPU under weak and strong scalability scenarios.
Figure 9 breaks down the dynamic energy consumption per DOF for the CG computation, with GPU and CPU contributions shown as colored segments in each bar (see legend). Under weak scalability, all implementations exhibit stable per-DOF energy consumption as the problem size increases. In other words, the energy required per DOF remains nearly constant when the number of processes grows proportionally to the global problem size, indicating that the additional computational resources are effectively amortized and that none of the implementations introduces significant energy overhead as the scale increases. The energy profile of the un-preconditioned CG solver follows the same relative trends observed for the SpMV operation in 5.1. This is expected, since SpMV is by far the dominant kernel within each iteration, and the improvements achieved by BootCMatchGX over Ginkgo in the standalone SpMV directly carry over to the full solver. In absolute terms, however, the dynamic energy per DOF in CG is about two orders of magnitude higher than in SpMV, reflecting the cumulative cost of performing many iterations rather than a single matrix–vector product. Although additional vector operations are executed in each iteration, their contribution does not alter the overall picture, which remains dictated by the efficiency of the SpMV kernel. This explains why BootCMatchGX (and, in the 7-point stencil case, NVIDIA AmgX) consistently achieve significantly lower dynamic energy consumption per iteration and per DOF compared to Ginkgo, as confirmed by the results in Figures 8 and 9. In summary, as expected, the comparative trends observed in CG largely reflect those already highlighted for SpMV, underlining the central role of the sparse matrix–vector product in defining the solver’s energy efficiency.
(a) 7-points stencil matrix with $408^{3}$ DOFs per GPU under weak scalability.
(b) 27-points stencil matrix with $265^{3}$ DOFs per GPU under weak scalability.
(c) 7-points stencil matrix with a total of $408^{3}$ DOFs under strong scalability.
(d) 27-points stencil matrix with a total of $265^{3}$ DOFs under strong scalability.
Figure 10: GPU power peak of the CG computation under weak and strong scalability scenarios.
Finally, Figure 10 shows the GPU power peaks recorded during the CG computation. BootCMatchGX consistently exhibits lower peaks than both Ginkgo and NVIDIA AmgX. Under strong scalability, these peaks decrease as the number of GPUs increases, reflecting the reduced workload per GPU. The higher power peaks observed for Ginkgo, in both weak and strong scalability, help explain its higher energy consumption even when, in the strong scaling scenario, execution times are comparable. Overall, this suggests that BootCMatchGX manages GPU resources more efficiently, maintaining a steady workload and avoiding sudden spikes in power demand.
5.3 Preconditioned Conjugate Gradient solver
We evaluate the PCG implementations of NVIDIA AmgX and BootCMatchGX under strong and weak scalability conditions. For fairness, AmgX is configured with the matching-based aggregation preconditioner, using aggregates of size 8, as in BootCMatchGX. Both libraries use default settings for the AMG hierarchy (levels and coarsest size) and apply the same smoother (4 $\ell_{1}$-Jacobi iterations) in the V-cycle. For the single-GPU case, the problem size is set to $370^{3}$ DOFs (7-point stencil). All performance measurements reported in the figures represent averages over five independent runs. For these tests, iterations stop at a relative residual tolerance of $10^{-6}$, reflecting realistic simulation settings.
(a) 7-points stencil matrix with $370^{3}$ DOFs pr GPU under weak scalability.
(b) 7-points stencil matrix with a total of $370^{3}$ DOFs under strong scalability.
Figure 11: Execution times breakdown of the PCG method of solve and setup times under weak and strong scalability scenarios. The number on top of each bar denotes the number of iterations carried out by solvers.
(a) 7-points stencil matrix with $370^{3}$ DOFs per GPU under weak scalability.
(b) 7-points stencil matrix with a total of $370^{3}$ DOFs under strong scalability.
Figure 12: Solve time per iteration of the PCG method under weak and strong scalability scenarios.
Figure 11 breaks down the PCG execution times in setup and solve phases shown as colored segments in each bar. The results indicate that BootCMatchGX consistently outperforms NVIDIA AmgX by achieving lower runtimes. This improvement stems from the BootCMatchGX preconditioner, which reduces the number of solver iterations and thus provides a clear reduction in the solution time. Moreover, as shown in Figure 12, the per-iteration solve time of BootCMatchGX also remains consistently lower than that of NVIDIA AmgX, thereby confirming the performance advantage of BootCMatchGX over NVIDIA AmgX.
(a) 7-points stencil matrix with $370^{3}$ DOFs per GPU under weak scalability.
(b) 7-points stencil matrix with a total of $370^{3}$ DOFs under strong scalability.
Figure 13: Dynamic energy consumption breakdown of the PCG computation on GPU and CPU under weak and strong scalability scenarios.
Figure 13 breaks down the dynamic energy consumption for the PCG computation, with GPU and CPU contributions shown as colored segments in each bar. The results confirm the trends observed in Figure 11: BootCMatchGX is consistently more energy‑efficient than NVIDIA AmgX. In every scenario, its dynamic energy consumption is lower than that of NVIDIA AmgX. As noted above, the lower iteration count of BootCMatchGX and per-iteration solve time yield an energy‑saving advantage over NVIDIA AmgX.
(a) 7-points stencil matrix with $370^{3}$ DOFs per GPU under weak scalability.
(b) 7-points stencil matrix with a total of $370^{3}$ DOFs under strong scalability.
Figure 14: Dynamic energy consumption per DOF breakdown of the PCG computation on GPU and CPU under weak and strong scalability scenarios.
Figure 14 presents the breakdown of dynamic energy consumption per DOF for the PCG computation, with GPU and CPU contributions shown as colored segments in each bar. Unlike the un-preconditioned CG case, and as expected given the additional cost of applying the preconditioner, we observe an increase in per-DOF energy consumption under both weak and strong scaling. Although the growth is gradual, it becomes noticeable as the number of tasks increases proportionally to the global problem size. The effect is more pronounced in the strong scaling scenario, where the reduced workload per GPU leads to lower computational efficiency. In all cases, however, BootCMatchGX remains consistently more energy-efficient than NVIDIA AmgX under both weak and strong scalability.
(a) 7-points stencil matrix with $370^{3}$ DOFs per GPU under weak scalability.
(b) 7-points stencil matrix with a total of $370^{3}$ DOFs under strong scalability.
Figure 15: Dynamic energy consumption per iteration breakdown of the PCG computation on GPU and CPU under weak and strong scalability scenarios.
Figure 15 presents the breakdown of dynamic energy consumption per iteration for the PCG computation, with GPU and CPU contributions shown as colored segments in each bar. Under weak scaling, the per-iteration energy increases moderately as the number of processes grows. This trend is consistent with the growing cost of communication/synchronization and the parallel overheads of applying the preconditioner at larger scales. Notably, the weak-scaling trends of BootCMatchGX and NVIDIA AmgX are essentially indistinguishable, indicating that the impact of the preconditioner application is equivalent for both implementations. In the strong-scaling scenario, as the workload per GPU diminishes, NVIDIA AmgX exhibits a slight advantage in energy efficiency over BootCMatchGX at higher process counts.
(a) 7-points stencil matrix with $370^{3}$ DOFs per GPU under weak scalability.
(b) 7-points stencil matrix with a total of $370^{3}$ DOFs under strong scalability.
Figure 16: GPU power peak of the PCG computation under weak and strong scalability scenarios.
Finally, Figure 16 shows the GPU power peaks recorded during the PCG computation. Under weak scalability, BootCMatchGX consistently exhibits lower peaks than NVIDIA AmgX, suggesting that it manages GPU resources more efficiently by maintaining a steady workload and avoiding sudden spikes in power demand. Under strong scalability, power peaks decrease as the number of GPUs increases, reflecting the reduced workload per GPU. However, starting from 8 GPUs, NVIDIA AmgX exhibits lower peaks than BootCMatchGX. This behavior is reflected in the higher per-iteration energy consumption of BootCMatchGX, as shown in Figure 15.
Overall, our analysis shows that BootCMatchGX outperforms NVIDIA AmgX in both execution time and energy consumption, thereby reducing the overall energy footprint. The only exception occurs in the strong scaling scenario, where NVIDIA AmgX demonstrates a slight advantage in per-iteration dynamic energy consumption as the number of GPUs increases.
6 Concluding Remarks and Future Work
This study demonstrates that optimizing numerical kernels and communication strategies in multi-GPU environments not only improves runtime performance, but also significantly reduces the energy footprint of large-scale sparse linear solvers. The results confirm that BootCMatchGX is competitive with, and in most cases superior to, established libraries such as Ginkgo and NVIDIA AmgX. Specifically, it achieves higher energy efficiency in SpMV and CG computations, while delivering a clear advantage in PCG thanks to its preconditioner design.
From an application perspective, these findings highlight the importance of considering energy efficiency as a key metric in the development of next-generation numerical libraries. Sustainable high-performance computing will increasingly depend on software that balances performance scalability with optimized energy use.
Future research will be directed toward:
  • Developing efficient and scalable AMG preconditioners that leverage mixed-precision arithmetic, with the goal of further reducing both execution time and energy consumption while preserving numerical robustness.
  • Validating the library’s energy benefits in multidisciplinary application domains, such as large-scale PDE simulations and scientific machine learning.
Acknowledgement
The authors would like to thank the developers of GPowerU for providing an effective tool for measuring GPU energy consumption, developed within the framework of the European TEXTAROSSA project and employed as the basis for this work. We also gratefully acknowledge Marcel Koch, from the Scientific Computing Center (SCC) at Karlsruhe Institute of Technology and developer of Ginkgo, for his support in its use.